### v1

@@ 3 Spectral energy distribution fitting
Physical properties of DESI galaxies are derived by performing SED (optical and mid-IR photometry) fitting using Code Investigating GALaxy Emission (CIGALE v2022.1; Boquien et al., 2019). CIGALE is a state-of-the-art Python code based on the principles of the energy balance between the dust-absorbed stellar emission in the ultraviolet (UV) and optical and its re-emission in the infrared (IR). Thanks to its efficiency, flexibility, and accuracy, CIGALE and its modified version X-CIGALE (Yang et al., 2020, 2022) are widely used to derive the physical properties of galaxies and AGN in large galaxy surveys (e.g. Ciesla et al. 2015; Salim et al. 2016, 2018; Małek et al. 2018; Barrows et al. 2021; Mountrichas et al. 2021b; Zou et al. 2022; Csizi et al. 2024; Osborne & Salim 2024) as well as high-z AGN (e.g. Conselice et al. 2023; Mezcua et al. 2023; Yang et al. 2023; Burke et al. 2024; Durodola et al. 2024). CIGALE estimates the physical properties of galaxies using a Bayesian approach by evaluating all the possible combinations of SED models on data to maximize the likelihood distribution. CIGALE takes into account the age of the universe at the redshift of each object when fitting models. It excludes stellar population ages that exceed the age of the universe at the given redshift. This constraint helps in avoiding unphysical solutions and ensures the consistency of the fitted parameters with cosmological constraints. The estimates and errors of the physical properties are then computed as the likelihood-weighted mean and standard deviation, respectively, for all the models (Boquien et al., 2019). To build a library of models, CIGALE relies on five main modules: star formation history (SFH), SSP models, dust attenuation and emission, and the AGN component. For each module, CIGALE includes several possible prescriptions and the flexible parametrization of model parameters allows us to adapt model complexity (i.e. number of free parameters). In the next Sections, we describe the key assumptions and parametrization of models used to create our catalog, which are summarized in Table 3. This configuration generates 167,529,600 models spanning over a wide redshift range from 0 to 6 (302,400 per redshift). We use a single node of 32 cores on the Cori supercomputer at the National Energy Research Scientific Computing Center (NERSC) to fit all SEDs within 3.5 hours.
@@ 3.1 Stellar component
We use the Bruzual & Charlot 2003 SSP models adopting a Chabrier (2003) IMF to build the stellar component. We assume solar metallicity and following the analysis by Ciesla et al. (2015), we use the delayed SFH with an optional exponential burst. This prescription allows us to reproduce the SEDs of both star-forming and passive galaxies with a modest number of free parameters (e.g. Ciesla et al., 2015; Salim et al., 2016). Such two exponentially decreasing star formation rate (SFR) laws with different e-folding times show a good performance in decoupling the long-term SFH from the recent star formation activity (Ciesla et al., 2015; Małek et al., 2018). The SFR is therefore defined as the sum of two exponentially decreasing SFRs:
@@
where t is the time, is the e-folding time of the main (old) stellar population, is the e-folding time of the burst (young) stellar population, and k is the amplitude of the second exponential which depends on the fraction of stars formed in the second burst versus the total stellar mass formed (). The SFH module is described in more detail in Ciesla et al. (2015, 2017).
The effects of the choice of IMF, the SFH prescription, and the solar metallicity assumption on the main galaxy physical properties are discussed in Sec. 6.
@@ 3.2 Nebular emission
Nebular emission (emission from ionized gas) is an important component to include when considering high-redshift galaxies (e.g. Stark et al., 2013; de Barros et al., 2014; Yuan et al., 2019) or at lower-redshift starburst dwarf and young star-forming galaxies (e.g Boquien et al., 2010) as intense star formation and ionization processes lead to stronger nebular emission signatures. Neglecting the nebular emission component may lead to the overestimation of the stellar mass (Yuan et al., 2019). The nebular emission lines are pre-computed with CLOUDY 17.01 (Ferland et al., 2017) with electron density (), gas metallicity (), and ionization parameter (U) as the free parameters and re-scaled with the number of Lyman continuum photons from the stellar emission. The Bruzual & Charlot (2003) SSP models using a constant SFH over 10 Myr is used to generate the photo-ionizing field shape. The nebular continuum is scaled directly from the number of ionizing photons. CIGALE takes into account also the fraction of the Lyman photons escaping galaxies and absorbed by dust. More details about the implementation of the CLOUDY into CIGALE can be found in Villa-Vélez et al. (2021). We keep the default parameters of this module (see Tab. 3).
@@ 3.3 Interstellar dust
Dust is a fundamental component of galaxies that significantly influences their observed SEDs, especially those that are actively star-forming (Conroy, 2013). It plays a dual role in stellar population models, as dust absorbs short-wavelength light (from the UV to NIR) and re-emits it at longer wavelengths (from MIR to far-infrared (FIR)). SED fitting techniques, such as those employed by CIGALE, model the full SED by simultaneously considering both the attenuated light and the re-emitted radiation. This integrated approach ensures that both the absorption and emission by dust are used to constrain the model. This is important because dust obscuration is influenced by the galaxy’s geometry, while dust emission is sensitive to the interstellar radiation field (Conroy 2013). Dust emission is dominated by: i) polycyclic aromatic hydrocarbon (PAH) bands in the MIR (), ii) very small, warm grains, and iii) big, relatively cold grains (). Differences between these dust grains have an impact on the dust SED. On the other hand, attenuation also depends on the geometry. The simplest way to model dust attenuation is to assume attenuation laws (Boquien et al., 2019). CIGALE provides two modules to model attenuation curves: the implementation of the Charlot & Fall (2000) model, and the modified Calzetti et al. (2000) model, we refer to it as the Calzetti et al. (2000) for simplicity.
The starburst model uses the Calzetti et al. (2000) starburst attenuation curve as a baseline, which is extended by a Leitherer et al. (2002) curve from the Lyman break to 150 nm. The amount of attenuation is described by the color excess applied to the nebular emission lines, , and the ratio , where is the color excess applied to the whole stellar continuum. Following the Calzetti et al. (2000) recommendations, this ratio is fixed to 0.44. We use the Calzetti et al. (2000) model to generate our catalog, but the impact on the physical property estimates when using Charlot & Fall (2000) is discussed in Sec. 6.2.
CIGALE provides five modules to describe the IR emission from dust: Casey (2012), Dale et al. (2014), Draine & Li (2007) and its updated version of the Draine et al. (2014) model, and Themis dust emission models from Jones et al. (2017). To create our catalog, we rely on the Draine et al. (2014) models, which account for very different physical conditions with a variety of radiation fields and a variable PAH emission providing a high flexibility. The model assumes that the majority () of dust mass is heated by a radiation field with an intensity (), while the remaining fraction () is exposed to intensities ranging from to following a power-law index . By default, and are fixed values set to 0.02, and 2, respectively. We also consider the model given by Dale et al. (2014), and describe its influence on the physical properties of galaxies in Sec. 6.2.
@@ 3.4 AGN contribution
CIGALE allows for the separation of the emission from AGN from their host galaxy with several approaches, starting from a simple AGN parameterization by the power slope when using Casey (2012) models to fit IR. The Dale et al. (2014) module provides simple templates of quasars from UV to IR, with a fraction of AGN as a free parameter. Those options are fast but do not provide complex AGN SEDs. However, CIGALE also incorporates two more flexible models: Fritz et al. (2006) and SKIRTOR (Stalevski et al., 2012, 2016). The AGN model from Fritz et al. (2006) covers the UV to IR and assumes that the central engine is surrounded by smoothly distributed dust in the AGN torus (i.e. the AGN unified model; e.g. Zou et al., 2019), while SKIRTOR models add the possibility that the dust is clumpy (e.g. Stalevski et al., 2012; Assef et al., 2013). However, it is still unclear whether observations can discriminate between these models (Feltre et al., 2012).
In this work, we apply the Fritz et al. (2006) model that uses a simple, but realistic torus geometry relying on the flared disc and a full range of dust grain size. It allows to control the geometry and physics of the torus by flexible ratio of the maximum to minimum radii of the dust torus, the optical depth at 9.7 , dust density distribution, opening angle of the torus, and the viewing angle. In particular, viewing from the equatorial direction (with a viewing angle AGNPSY ) leads to the obscuration of the central engine and only the radiation reemitted in IR can be observed (type 2 AGN, i.e. NL AGN). When viewing from the polar direction (with a viewing angle AGNPSY ), the central engine is directly visible (type 1 AGN, i.e. BL AGN). We fix the parameters to default, except allowing for the flexibility in the viewing angle (AGNPSY) and the AGN fraction (AGNFRAC) defined as the ratio of the AGN IR emission to the total IR emission. Thus, the generated models cover a wide range of objects, including galaxies without AGN contribution, as well as type 1 and 2 AGN. This provides flexibility and allows us to build the catalog of physical properties for both AGN and non-AGN host galaxies. The impact of the incorporation or the change of the AGN model is discussed in Sec. 6.3.
@@ 4 VAC: general properties
In this Section, we characterize the VAC of physical properties of the DESI galaxies and quasars observed at redshift (see Sec. 2 for data description). The catalog includes all DESI sources, independent of their target class (BGS, LRG, ELG, SCND, MWS and QSO) or redshift. The physical properties are derived based on the SED fitting using CIGALE (see Sec. 3 for the description of the SED procedures and modules and Table 3 for a full list of parameters). The catalog includes estimates of stellar masses, SFRs, absolute magnitudes, AGN fractions, and AGN luminosities among others (for the full list see Table 8).
4.1 Selection of galaxies - SEDs with secure fits
@@
The DESI VAC includes all ‘reliable’ sources from the EDR (see Sec. 2 for the cuts applied for sample selection). The choice of additional cleaning cuts strongly depends on individual scientific goals. For general analysis of DESI galaxies, we do not introduce strict cuts to clean the sample, but the cuts can be adjusted as required for the specific science case. In the following analysis, we exclude:
@@
153,294 (12%) sources for which . Almost all sources with are characterized by insufficient photometric observations to perform a reliable SED fitting. In particular, half of them are observed in only one optical band with a high signal to noise ratio (SNR ); 21% of them have two or three optical observations with SNR , and only 3% are observed in two or more WISE bands with SNR ). This suggests that these sources are false detections, faint sources, and other artifacts without valid photometry.
@@
An additional 3,231 (0.2%) sources with bad fits characterized by (see. App. D.2 for the description of the cut). We note that the user may apply a more restrictive cut on (e.g., ) or apply further statistical criteria, such as a Bayesian information criterion (see e.g. Masoura et al. 2018; Buat et al. 2021). Figure 2 shows the distribution of for DESI EDR galaxies.
@@
Additionally, the user can introduce several other quality cuts based on additional flags provided in the catalog. Namely:
@@
SNR of the optical photometry. The uncertainties in stellar mass estimates strongly depend on the input photometry (see App. B for details). Sources observed in three optical bands (grz) with high SNR (SNR ) are characterized by a stellar mass error of , corresponding to the standard uncertainty of the stellar mass estimates due to model assumptions (see Sec. 6 and Conroy 2013777Conroy (2013) showed that different assumptions of the IMF, SSP models, SFH can introduce a systematic uncertainty of dex., Pacifici et al. 2023). The FLAGOPTICAL defines the number of optical bands with SNR and can be used to select sources with more reliable photometry and thus physical properties estimates.
@@
SNR of the WISE photometry. The availability of WISE photometry with high SNR (SNR ) has an impact not only on the stellar mass estimates (see Sec. 6.4) but also on stellar mass errors (see App. B). The FLAGINFRARED defines the number of bands (W1-4) with SNR 888The catalog includes the SNR for each band allowing the users to modify the threshold according to their scientific cases..
@@
The probability density function (PDF) of the estimated parameters is asymmetric or multi-peaked. CIGALE introduces two estimates based on the best-fit model (best) and the likelihood-weighted mean measured from its PDF marginalized over other parameters (bayes). A narrow, one-peak PDF should have these two values very similar, otherwise the PDF is asymmetric or multi-peaked. The shape of the PDF for the stellar mass and SFR estimates is expressed as and , respectively. To preselect sources with narrow one-peak PDF of stellar mass (SFR) one can consider only the one with values between following Mountrichas et al. (2021b), and Mountrichas et al. (2024).
@@
Additional cuts based on the Tractor photometry information. To reject fragmented sources one may introduce the cut: FRACFLUX as advised by Juneau et al. (2024), Pucha et al. under DESI Collaboration review.
@@
Our final EDR sample after cleaning includes 1,129,599 sources (88% of the whole catalog) and is characterized by a mean (see Fig. 2). At least of the sample is characterized by good fits defined with a more strict criterion of (Masoura et al., 2018; Buat et al., 2021).
@@
The distributions of SED-derived properties: rest-framed g-r color, stellar mass, and SFR of the DESI EDR galaxies are shown in Fig. 3. The distribution shapes are clearly different for each of the main target classes (MWS is scaled up by 10 to ease the comparison). The LRG are found among the reddest observed galaxies, while ELG are among the bluest with BGS bridging both the blue and red ends of the distribution (see the top panel in Fig. 3). QSO, MWS, and SCND are among the blue population. The stellar mass distribution follows the distribution of the g-r color, i.e. the redder target class, LRG, is also found to be the one covering the high-mass end of the DESI EDR distribution (), while the blue ELG peaks at forming a long tail towards the low-mass end. The remaining target classes (BGS, QSO, SCND, and MWS) are peaking in-between ELG and LRG covering . As expected, LRG are characterized by lower SFR () while ELG peak at higher SFR () with MWS found to be among the most active ().

### v2

@@ 3 Spectral energy distribution fitting
Physical properties of DESI galaxies are derived by performing SED (optical and mid-IR photometry) fitting using Code Investigating GALaxy Emission (CIGALE v2022.1; Boquien et al., 2019). CIGALE is a state-of-the-art Python code based on the principles of the energy balance between the dust-absorbed stellar emission in the ultraviolet (UV) and optical and its re-emission in the infrared (IR). Thanks to its efficiency, flexibility, and accuracy, CIGALE and its modified version X-CIGALE (Yang et al., 2020, 2022) are widely used to derive the physical properties of galaxies and AGNs in large galaxy surveys (e.g., Ciesla et al. 2015; Salim et al. 2016, 2018; Małek et al. 2018; Barrows et al. 2021; Mountrichas et al. 2021b; Zou et al. 2022; Csizi et al. 2024; Osborne & Salim 2024) as well as high-z AGNs (e.g., Conselice et al. 2023; Mezcua et al. 2023; Yang et al. 2023; Burke et al. 2024; Durodola et al. 2024). CIGALE estimates the physical properties of galaxies using a Bayesian approach by evaluating all the possible combinations of SED models on data to maximize the likelihood distribution. CIGALE takes into account the age of the universe at the redshift of each object when fitting models. It excludes stellar population ages that exceed the age of the universe at the given redshift. This constraint helps us avoid unphysical solutions and ensures the consistency of the fitted parameters with cosmological constraints. The estimates and errors of the physical properties are then computed as the likelihood-weighted mean and standard deviation, respectively, for all the models (Boquien et al., 2019). To build a library of models, CIGALE relies on five main modules: star formation history (SFH), SSP models, dust attenuation and emission, and the AGN component. For each module, CIGALE includes several possible prescriptions and the flexible parametrization of model parameters allows us to adapt the model complexity (i.e., number of free parameters). In the next sections, we describe the key assumptions and parametrization of models used to create our catalog, which are summarized in Table 3. This configuration generates 167,529,600 models spanning over a wide redshift range from 0 to 6 (302,400 per redshift). We use a single node of 32 cores on the Cori supercomputer at the National Energy Research Scientific Computing Center (NERSC) to fit all SEDs within 3.5 hours.
@@ 3.1 Stellar component
We use the Bruzual & Charlot 2003 SSP models adopting a Chabrier (2003) IMF to build the stellar component. We assume solar metallicity and following the analysis by Ciesla et al. (2015), we use the delayed SFH with an optional exponential burst. This prescription allows us to reproduce the SEDs of both star-forming and passive galaxies with a modest number of free parameters (e.g., Ciesla et al., 2015; Salim et al., 2016). Such two exponentially decreasing star formation rate (SFR) laws with different e-folding times show a good performance in decoupling the long-term SFH from the recent star formation activity (Ciesla et al., 2015; Małek et al., 2018). The SFR is therefore defined as the sum of two exponentially decreasing SFRs:
@@
where t is the time since the onset of star formation, is the e-folding time of the main (old) stellar population, is the e-folding time of the burst (young) stellar population, and k is the normalization constant which depends on the fraction of stars formed in the second burst versus the total stellar mass formed (). The SFH module is described in more detail in Ciesla et al. (2015, 2017).
The effects of the choice of IMF, the SFH prescription, and the solar metallicity assumption on the main galaxy physical properties are discussed in Sect. 6.
@@ 3.2 Nebular emission
Nebular emission (emission from ionized gas) is an important component to include when considering high-redshift galaxies (e.g., Stark et al., 2013; de Barros et al., 2014; Yuan et al., 2019) or at lower-redshift starburst dwarf and young star-forming galaxies (e.g., Boquien et al., 2010) as intense star formation and ionization processes lead to stronger nebular emission signatures. Neglecting the nebular emission component may lead to the overestimation of the stellar mass (Yuan et al., 2019). The nebular emission lines are pre-computed with CLOUDY 17.01 (Ferland et al., 2017) with electron density (), gas metallicity (), and ionization parameter (U) as the free parameters and re-scaled with the number of Lyman continuum photons from the stellar emission. The Bruzual & Charlot (2003) SSP models using a constant SFH over 10 Myr is used to generate the photo-ionizing field shape. The nebular continuum is scaled directly from the number of ionizing photons. CIGALE takes into account also the fraction of the Lyman photons escaping galaxies and absorbed by dust. More details about the implementation of the CLOUDY into CIGALE can be found in Villa-Vélez et al. (2021). We retained the default parameters of this module (see Table 3).
@@ 3.3 Interstellar dust
Dust is a fundamental component of galaxies that significantly influences their observed SEDs, especially those that are actively star-forming (Conroy, 2013). It plays a dual role in stellar population models, as dust absorbs short-wavelength light (from the UV to NIR) and re-emits it at longer wavelengths (from MIR to far-infrared (FIR)). SED fitting techniques, such as those employed by CIGALE, model the full SED by simultaneously considering both the attenuated light and the re-emitted radiation. This integrated approach ensures that both the absorption and emission by dust are used to constrain the model. This is important because dust obscuration is influenced by the galaxy’s geometry, while dust emission is sensitive to the interstellar radiation field (Conroy 2013). Dust emission is dominated by: i) polycyclic aromatic hydrocarbon (PAH) bands in the MIR (), ii) very small, warm grains, and iii) big, relatively cold grains (). Differences between these dust grains have an impact on the dust SED. On the other hand, attenuation also depends on the geometry. The simplest way to model dust attenuation is to assume attenuation laws (Boquien et al., 2019). CIGALE provides two modules to model attenuation curves: the implementation of the Charlot & Fall (2000) model and the modified Calzetti et al. (2000) model; for simplicity, we refer to it as the Calzetti et al. (2000) model.
The starburst model uses the Calzetti et al. (2000) starburst attenuation curve as a baseline, which is extended by a Leitherer et al. (2002) curve from the Lyman break to 150 nm. The amount of attenuation is described by the color excess applied to the nebular emission lines, , and the ratio , where is the color excess applied to the whole stellar continuum. Following the Calzetti et al. (2000) recommendations, this ratio is fixed to 0.44. We used the Calzetti et al. (2000) model to generate our catalog; however, the impact on the physical property estimates when using Charlot & Fall (2000) is discussed in Sect. 6.2.
CIGALE provides five modules to describe the IR emission from dust: Casey (2012), Dale et al. (2014), Draine & Li (2007) and its updated version of the Draine et al. (2014) model, along with the Themis dust emission models from Jones et al. (2017). To create our catalog, we rely on the Draine et al. (2014) models, which account for very different physical conditions with a variety of radiation fields and a variable PAH emission providing a high flexibility. The model assumes that the majority () of dust mass is heated by a radiation field with an intensity (), while the remaining fraction () is exposed to intensities ranging from to following a power-law index . By default, and are fixed values set to 0.02, and 2, respectively. We also considered the model from Dale et al. (2014) and describe its influence on the physical properties of galaxies in Sect. 6.2.
@@ 3.4 AGN contribution
CIGALE allows for the separation of the emission from AGN from their host galaxy with several approaches, starting from a simple AGN parameterization by the power slope when using Casey (2012) models to fit IR data. The Dale et al. (2014) module provides simple templates of quasars from UV to IR, with the AGN fraction left as a free parameter. Those options are fast but do not provide complex AGN SEDs. However, CIGALE also incorporates two more flexible models: Fritz et al. (2006) and SKIRTOR (Stalevski et al., 2012, 2016). The AGN model from Fritz et al. (2006) covers the UV to IR and assumes that the central engine is surrounded by smoothly distributed dust in the AGN torus (i.e., the AGN unified model; e.g., Zou et al., 2019), while SKIRTOR models add the possibility that the dust is clumpy (e.g., Stalevski et al., 2012; Assef et al., 2013). However, it is still unclear whether observations can discriminate between these models (Feltre et al., 2012).
In this work, we apply the Fritz et al. (2006) model, which uses a simple, but realistic torus geometry that relies on the flared disc and a full range of dust grain size. It allows us to control the geometry and physics of the torus thanks to a flexible ratio of the maximum to minimum radii of the dust torus, the optical depth at 9.7 , dust density distribution, opening angle of the torus, and the viewing angle. In particular, viewing from the equatorial direction (with a viewing angle AGNPSY ) leads to the obscuration of the central engine and only the radiation reemitted in IR can be observed (type 2 AGN, i.e., NL AGN). When viewing from the polar direction (with a viewing angle AGNPSY ), the central engine is directly visible (type 1 AGN, i.e., BL AGN). We fixed the parameters to the default, aside from allowing for the flexibility in the viewing angle (AGNPSY) and in the AGN fraction (AGNFRAC), defined as the ratio of the AGN IR emission to the total IR emission. Thus, the generated models cover a wide range of objects, including galaxies without any AGN contribution, as well as type 1 and 2 AGNs. This provides flexibility and allows us to build the catalog of physical properties for both AGN and non-AGN host galaxies. The impact of the incorporation or the change of the AGN model is discussed in Sect. 6.3.
@@ 4 VAC: General properties
In this section, we characterize the VAC of physical properties of the DESI galaxies and quasars observed at redshift (see Sect. 2 for data description). The catalog includes all DESI sources, independent of their target class (BGS, LRG, ELG, SCND, MWS, and QSO) or redshift. The physical properties are derived based on the SED fitting using CIGALE (see Sect. 3 for the description of the SED procedures and modules and Table 3 for a full list of parameters). The catalog includes estimates of stellar masses, SFRs, absolute magnitudes, AGN fractions, and AGN luminosities among others (for the full list, see Table 8).
4.1 Selection of galaxies: SEDs with secure fits
@@
The DESI VAC includes all ‘reliable’ sources from the EDR (see Sect. 2 for the cuts applied for sample selection). The choice of additional cleaning cuts strongly depends on individual scientific goals. For general analysis of DESI galaxies, we did not introduce strict cuts to clean the sample, but the cuts can be adjusted as required for the specific science case. In the following analysis, we have excluded:
@@
153,294 (12%) sources for which . Almost all sources with are characterized by insufficient photometric observations to perform a reliable SED fitting. In particular, half of them are observed in only one optical band with a high signal to noise ratio (S/N ); 21% of them have two or three optical observations with S/N , and only 3% are observed in two or more WISE bands with S/N ). This suggests that these sources are false detections, faint sources, and other artifacts without valid photometry.
@@
An additional 3,231 (0.2%) sources with bad fits characterized by (see Appendix D.2 for the description of the cut). We note that the user may apply a more restrictive cut on (e.g., ) or apply further statistical criteria, such as a Bayesian information criterion (BIC; see e.g., Masoura et al. 2018; Buat et al. 2021). Figure 2 shows the distribution of for DESI EDR galaxies.
@@
Additionally, the user can introduce several other quality cuts based on additional flags provided in the catalog, namely:
@@
The S/N of the optical photometry. The uncertainties in stellar mass estimates strongly depend on the input photometry (see Appendix B for details). Sources observed in three optical bands (grz) with high S/N (S/N ) are characterized by a stellar mass error of , corresponding to the standard uncertainty of the stellar mass estimates due to model assumptions (see Sect. 6 and Conroy 2013999Conroy (2013) showed that different assumptions of the IMF, SSP models, SFH can introduce a systematic uncertainty of dex., Pacifici et al. 2023). The FLAGOPTICAL defines the number of optical bands with S/N and can be used to select sources with more reliable photometry and, thus, better estimates of the physical properties.
@@
The S/N of the WISE photometry. The availability of WISE photometry with high S/N (S/N ) has an impact not only on the stellar mass estimates (see Sect. 6.4), but also on stellar mass errors (see App. B). The FLAGINFRARED defines the number of bands (W1-4) with S/N 101010The catalog includes the S/N for each band allowing the users to modify the threshold according to their scientific cases..
@@
The probability density function (PDF) of the estimated parameters is asymmetric or multi-peaked. CIGALE introduces two estimates based on the best-fit model (best) and the likelihood-weighted mean measured from its PDF marginalized over other parameters (bayes). A narrow, one-peak PDF should see very similar values for these estimates; otherwise the PDF would end up asymmetric or multi-peaked. The shape of the PDF for the stellar mass and SFR estimates is expressed as and , respectively. To preselect sources with narrow one-peak PDF of stellar mass (SFR), we can consider only the one with values between following Mountrichas et al. (2021b), and Mountrichas et al. (2024).
@@
Additional cuts based on the Tractor photometry information. To reject fragmented sources, we may introduce the cut: FRACFLUX as advised by Juneau et al. (2024), Pucha et al. under DESI Collaboration review.
@@
Our final EDR sample after cleaning features 1,129,599 sources (88% of the whole catalog) and is characterized by a mean (see Fig. 2). At least of the sample is characterized by good fits defined with a more strict criterion of (Masoura et al., 2018; Buat et al., 2021).
@@
The distributions of SED-derived properties: rest-frame g-r color, stellar mass, and SFR of the DESI EDR galaxies are shown in Fig. 3. The distribution shapes are clearly different for each of the main target classes (MWS is scaled up by 10 to ease the comparison). The LRG are found among the reddest observed galaxies, while ELG are among the bluest with BGS bridging both the blue and red ends of the distribution (see the top panel in Fig. 3). QSO, MWS, and SCND are among the blue population. The stellar mass distribution follows the distribution of the g-r color, namely, the redder target class, LRG, is also found to be the one covering the high-mass end of the DESI EDR distribution (), while the blue ELG peaks at forming a long tail towards the low-mass end. The remaining target classes (BGS, QSO, SCND, and MWS) are peaking in-between ELG and LRG covering . As expected, LRG are characterized by lower SFR () while ELG peak at higher SFR (), with MWS found to be among the most active ().